# Feedback Connections in Quantum Reservoir Computing with Mid-Circuit Measurements

### v1

@@ I Introduction
Quantum computers have proven to provide significant speedups for specialized problems by exploiting the unique principles of quantum mechanics. These improvements are also thought to be valuable for machine learning tasks [1, 2, 3, 4], although concrete evidence for an advantage is still a highly debated topic [5]. Quantum Reservoir Computing (QRC) is a promising approach in quantum machine learning with current noisy intermediate scale quantum hardware, due to an architecture that does not require expensive circuit optimization.
@@
In particular, reservoir computing is a machine learning approach for predicting time series data [6, 7, 8, 9]. A reservoir computer transforms an input sequence into a higher dimensional feature space by the use of a fixed but random non-linear system called a reservoir. A simple linear regression model is then trained to read the state of the reservoir and map the result to a desired output. Since the training is done only in the last stage, while the internal dynamics are fixed, the optimization is very efficient. QRC is the quantum analog of classical reservoir computing, where a quantum system is used as a reservoir and its complex dynamics are used to efficiently process temporal data [10, 11, 12, 13, 14, 15, 16]. This is seen as a promising approach because quantum systems offer a naturally high-dimensional Hilbert space and complex non-linear dynamics. These properties are thought to enhance the performance of reservoir computing [17, 14, 18].
@@
Several architectures have been proposed, broadly categorized into restart-based and continuous methods. Conventional restart-based QRC schemes generate the output time series by averaging the measurement results across all qubits at each time step. This requires repeatedly re-preparing and running the system from the beginning for each time step to obtain the complete time series, see Fig. 1 (a). To avoid rerunning the entire process from the beginning, methods known as rewind protocols have been proposed [15]. Here, the reservoir for cycle $k$ is initiated from cycle $k+1-\tau$ therefore reducing computation to the last $\tau$ cycles, as compared to the first cycle in the conventional approach (Fig. 1 (b)). This method is motivated by the fact reservoirs have a fading-memory property [19], which means that over time the reservoir loses information from distant timesteps, therefore making them insignificant for newer cycles. On the other hand, continuous QRC methods leverage mid-circuit measurements to continuously monitor the dynamics of the reservoir, Fig. 1 (d), improving overall execution time on quantum hardware immensely [20]. This method has also been combined with weak measurements to mitigate the erasure of past inputs, as they extract limited information without fully collapsing the quantum state. However, this comes at a cost, as the readout from the reservoir is now incomplete [15].
@@
A novel approach introduces a feedback link by feeding back expectation values into the reservoir [14]. This approach still restarts for each time step, but only needs to run for one time step at a time. While this method effectively incorporates memories of previous inputs into the reservoir, it still has a drawback, namely execution time on quantum hardware. As shown in Fig. 1 (d), for each time step the expectation value must be estimated with a certain number of shots.
@@
In this paper we introduce a novel feedback driven quantum reservoir computing approach with mid circuit measurements and classical feedforward operations during the coherence time of a qubit. As seen in Fig. 1 (d), each cycle requires $n$ shots. Thus, for an input sequence with a training length $t$ tested over $k$ different values, the total number of shots scales with $t\times k\times n$. As $t$, $k$, or preferably $n$ grow, the cumulative cost of running these circuits becomes large. This motivates our exploration of feedback connections that operate within a qubit’s coherence time, using novel classical feedforward methods. To isolate and rigorously quantify the effect of the mid-circuit feedback, we perform a qubit reset after each time step. This ensures no extraneous information (other than the intended feedback) propagates forward, allowing us to examine the feedback connection itself in a controlled manner. One may omit these resets to explore more complex scenarios in which correlations persist over time, but here our primary goal is to study feedback connections in isolation, thereby clarifying their impact. For implementation, IBM superconducting systems have notably demonstrated real-time classical feedforward operations [21], offering an option to reduce the overall shot overhead for certain problems. We use the measurement strings obtained through a mid-circuit measurement as feedback for the next cycle. We can tune the feedback strength and the input strength through hyperparameters. We evaluate the memory performance through investigation of the short-term memory task and evaluate the predictive capabilities with the prediction of the Mackey-Glass system [22] and the expectation value of one spin in a chaotic Ising chain.
@@
Figure 1: Quantum reservoir computing protocols: Each architecture is designed to process an input sequence ${s_{k}}$. The systems are initialized in the state $\rho_{0}=\ket{0}\bra{0}$ |. A unitary operator $U(\cdot)$ constructed from parametrized Pauli rotations, governs the state evolution. Each $U(\cdot)$ represents the processing of one timestep (cycle). (a) Graphical representation of a conventional restart based QRC regime. For each timestep the architecture needs to be restarted and run from the beginning. This amounts to a quadratic runtime. b A rewind protocol, that is not restarted from the beginning, but only processes the last $\tau$ timesteps. c A feedback driven QRC protocol. Only one timestep is processed at a time, but the expectation value from the previous experiment is put into reservoir. d A mid-circuit measurement architecture, that continously monitors the quantum state. No restarting is needed. This architecture can be combined with weak measurements. d A feedback driven QRC architecture with mid-circuit measurements. Feedback is incorporated by feeding the measurement strings $\textbf{m}_{k}$ back into the reservoir, a process that can be performed at runtime.
@@ II-A Architecture
We propose a QRC architecture with mid-circuit measurements, that incorporates feedback connections, as depicted in Fig. 1 (e). We start with a $N$ qubit quantum system, initialized in the state $\rho_{0}^{N}=\ket{0}\bra{0}^{\otimes N}$. Given an input sequence ${s_{k}}$, the system evolution is governed by the unitary operator $U(s_{k},\textbf{m}_{k-1})$, where $s_{k}$ is the current timestep of the input sequence and $\textbf{m}_{k-1}\in\{-1,1\}^{N}$ is previous measurement results of cycle $s_{k-1}$. The input sequence is placed into a two qubit gate $R(a_{in}s_{k})$, whereas each previous measurement result $m_{k-1}^{j}$ is inserted into another two qubit gate $R(a_{fb}m_{i}^{j})$. The structure of the $R$ gate is defined as:
@@
Here, $\theta$ represents an angle parameter that determines the rotation applied in the two-qubit gate. Specifically, $\theta$ can take values from the set $\{a_{in}s_{k},a_{fb}m_{k-1}^{j}\}$ where $a_{fb}$ and $a_{fb}$ are scaling parameters in $\mathcal{R}$, that allow to control the input and feedback strength. We want to add that this structure is an usual reservoir structure found in QRC literature [14, 20, 23]. The full unitary $U(s_{k},\textbf{m}_{k-1})$ consists of one input $R$ gate, $N$ feedback $R$ gates and one Haar random unitary. The idea is illustrated for the $2$ qubit case in Fig. 2. For the first timestep $s_{0}$ a random vector $\textbf{m}_{-1}$ uniformly distributed from $\{-1,1\}^{N}$ is selected for timestep 00. After unitary evolution the system in projectively measured in computational basis. We reset our system to $\rho_{0}^{N}$ and feed the measurement string forward into the next unitary.
@@ II-B Training
The task of quantum reservoir computing is defined by an input sequence $\{s_{k}\}$ and a target sequence $\{\hat{y}_{k}\}$. The goal is to build a regression model that accurately approximates the target sequence. In typical quantum reservoir computing mechanisms, there are three main stages: washout, training, and testing. As discussed in Section II-A, we begin the first cycle $s_{0}$ with a randomly selected feedback vector. To remove the effects of this initialization, we discard the first $l_{w}$ measurement outcomes (the washout phase). Subsequently, the following $l_{\mathrm{tr}}$ steps are used to train the model (the training phase), and performance is then evaluated on the next $l_{\mathrm{ts}}$ steps (the testing phase). The linear regression model is then build based on the following reservoir outputs. Specifically we build the matrix $X_{tr}\in\mathcal{R}^{l_{tr}\times(N+1)}$ defined as:
@@
where $\langle Z\rangle_{k}^{n}$ is the expectation value of the Pauli $Z$ operator on the nth qubit at the kth timestep.
We create now a linear model $w_{opt}$ such that $y_{tr}=X_{tr}w_{opt}$ closely resembles the target sequence $\hat{y}_{tr}$. We utilize Moore-Penrose pseudoinverse:
@@ II-C Testing
The testing routine largely follows the training routine; however, instead of creating a new model, we use the existing model to obtain a prediction vector of length $l_{ts}$. We create a matrix $X_{ts}\in\mathcal{R}^{l_{ts}\times(N+1)}$ defined as:
@@
The resulting prediction $y_{pred}$ is then compared against the target vector $\hat{y}_{ts}$ to assess the model’s predictive performance.
Figure 2: Minimal two-qubit unitary circuit. The circuit processes the input $s_{k}$ through a rotation $R(a_{\mathrm{in}}\,s_{k})$ and then incorporates feedback via rotations $R(a_{\mathrm{fb}}\,m_{k-1}^{0})$ and $R(a_{\mathrm{fb}}\,m_{k-1}^{1})$, where $\{m_{k-1}^{0},m_{k-1}^{1}\}$ are the prior measurement outcomes. The structure of $R(\cdot)$ is explained in Eq. 1. A final Haar-random unitary $U_{\mathrm{Haar}}$ is applied at the end.
@@ III-A Short-term memory capacity
To characterize the performance of remembering past inputs we use the short-term memory capacity [24]. Consider an input sequence $\{s_{k}\}$, where each $s_{k}$ is drawn independently from a uniform distribution on $[0,1]$. The objective is to retrieve past values of the sequence, formulated as $\hat{y}_{k}=s_{k+\tau}$, where $\tau$ is zero or negative. This metrics then utilizes the square of the correlation coefficient between the predicted target and real target:
$R^{2}_{\tau}=\frac{\text{cov}^{2}(\hat{y}_{ts},y_{pred})}{\sigma^{2}(\hat{y}_{ts})\sigma^{2}(y_{pred})}$ (6)
When $R^{2}_{\tau}$ is near 1, the model’s predictions align closely with the actual target values, indicating minimal error. In contrast, when $R^{2}_{\tau}$ is close to 0, the predicted outputs show substantial divergence from the true targets. The short-term memory capacity is then measured by summing up $R^{2}$ for various steps:
@@ III-B Predictive Capabilities
To evaluate predictive capabilities we test the model on one classical task and one quantum time series. For the study of classical dynamical systems, we select the Mackey-Glass time series, a well-known benchmark that has been extensively in the context of reservoir computing [25, 14, 10, 26]. The delay differential equation is defined as [22]:
@@
We set $\beta_{0}=0.2$, $\theta=1$, $n=10$, $\tau=17$ and $\gamma=0.1$. To generate the time series, we numerically integrate the delay differential equation using a forward Euler scheme with a time step $\Delta t=1$. The initial values are set to a constant for the duration of the delay, and the trajectory is normalized to the interval [0,1].
The quantum time series is constructed from the expectation value of the middle spin’s $\sigma^{z}$ operator in a one-dimensional Ising chain governed by the Hamiltonian
@@
We simulate the system using this Hamiltonian with parameters set to $N=5$, $J=1$, $h_{X}=1.05$, and $h_{Z}=-0.5$. At this parameter set the quantum Ising model becomes non integrable and exhibits chaotic behavior [27]. At each timestep, the expectation value $\langle Z^{2}_{t}\rangle$ of the central spin is measured, generating the time series. The time evolution is discretized with a timestep of $\Delta t=0.005$.
For all predictive tasks the goal is to forecast steps in the future, similarly formulated with the target sequence as $\hat{y}_{k}=s_{k+\tau}$, where $\tau$ is positive. We evaluate the models performance with the norm mean squared error, which is defined as:
@@
The NMSE measures the size of the error relative to the scale (magnitude) of the true data. A low error indicates a good performance of the model, whereas high values indicate a large deviation from the target sequence.
@@ III-C Analysis of Echo State Property
As previously mentioned in section I, one of the important aspects of a reservoir computing model is the so called fading-memory property. In the classical setting of echo state network (ESN) reservoir computing models, this can be implemented by assuring that the ESN contains the echo state property (ESP) [28], which ensures that the internal state of the model becomes asymptotically dependent on only the input history regardless of its initial conditions. However in quantum reservoir computing models, the fading-memory property is usually generated by including some (preferably controlled) dissipation into the model[29, 30]. Therefore the importance of the ESP in quantum reservoir models is still unclear.
As a first step into investigating the role of the ESP in quantum reservoir computing models, we perform an analysis of the asymptotic dependency of the reservoir state of several reservoir computing models (including the one proposed in this paper) with respect to their initial internal states. To do this, we look at how the internal state of the reservoir evolves over time across different scenarios of random initializations of the reservoir state. We can define the internal state of the reservoir $x_{t}\in\mathcal{R}^{M}$ at time $t$ to be:
@@
where $M$ is the reservoir dimension, $F$ is the fixed reservoir function and $s_{t}$ is the time series at time $t$ [28, 19]. [^1] In our architecture $x_{t}$ = $[\langle Z\rangle^{1}_{t},\langle Z\rangle^{2}_{t},\cdots,\langle Z\rangle^{N}_{t}]$ For each model, we use input sequences drawn from a uniform distribution on $[0,1]$ (as described in Section III-A) across 5 random initializations. For each run, we extract the first component of the reservoir’s internal state, $x_{t}^{1}$, at time $t$ and track its evolution over time across the different runs. To quantify the asymptotic behavior of these internal states, we compute the average difference of the feature across runs. We perform this analysis for a classical minimal ESN model [31], a feedback-driven QRC protocol [14], a QRC scheme with a mid-circuit measurement architecture [20], and the model proposed in this paper.

### v3

@@ I Introduction
Reservoir computing is a machine learning approach for predicting time series data [1, 2, 3, 4]. A reservoir computer transforms an input sequence into a higher dimensional feature space using a fixed but random non-linear system called a reservoir. A simple linear regression model is then trained to read the state of the reservoir and map the result to a desired output. Since the training is done only in the last stage, while the internal dynamics are fixed, the optimization is very efficient. Quantum Reservoir Computing (QRC) is the quantum analog of classical reservoir computing, where a quantum system is used as a reservoir and its complex dynamics are used to efficiently process temporal data [5, 6, 7, 8, 9, 10, 11]. This is seen as a promising approach because quantum systems offer a naturally high-dimensional Hilbert space and complex non-linear dynamics. These properties are thought to enhance the performance of reservoir computing [12, 9, 13].
@@
Several architectures have been proposed, broadly categorized into restart-based and continuous methods. Conventional restart-based QRC schemes generate the output time series by averaging the measurement results over all qubits at each time step. This requires repeatedly restarting the system for each time step to obtain the complete time series, see Fig. 1 (a). To avoid running the whole process from the beginning, methods known as rewind protocols have been proposed [10]. Here, the reservoir for cycle $k$ is initiated from cycle $k+1-\tau$, thus reducing the computation to the last $\tau$ cycles compared to the first cycle in the conventional approach (Fig.  1 (b)). This method is motivated by the fact that reservoirs have a fading memory property [14], which means that over time the reservoir loses information from distant time steps, making them insignificant for more recent cycles. On the other hand, continuous QRC methods use mid-circuit measurements to continuously monitor the dynamics of the reservoir, Fig. 1(d), which immensely improves the overall execution time on quantum hardware [15]. This method has also been combined with weak measurements to mitigate the erasure of past inputs, as they extract limited information without completely collapsing the quantum state. However, this comes at a cost, as the readout from the reservoir is now incomplete [10].
@@
A novel approach introduces a feedback link by feeding back expectation values into the reservoir [9]. This approach still restarts at each time step, but only needs to run for one time step at a time. While this method effectively incorporates memories of previous inputs into the reservoir, it still has a drawback, namely execution time on quantum hardware. As shown in Fig. 1 (c), for each time step the expectation value has to be estimated with a certain number of shots. Recently, a preprint has been published describing a modification of this protocol using weak measurements, where part of the reservoir state is carried over between time steps [16]. However, this method still requires many measurements to be taken between time steps in order to compute expectation values before feeding them into the next time step.
@@
In this paper, we introduce a novel feedback-driven quantum reservoir computing approach with mid-circuit measurements and classical feedforward operations during the coherence time of a qubit. In the restart based feedback protocol, as shown in Fig. 1 (c), each cycle requires $n$ shots. Thus, for an input sequence of training length $t$ tested over $k$ different reservoir realizations, the total number of shots scales with $t\times k\times n$. As $t$, $k$, or preferably $n$ grow, the cumulative cost of running these circuits becomes large. This motivates our exploration of feedback circuits operating within the coherence time of a qubit using novel classical feedforward methods. To isolate and rigorously quantify the effect of mid-circuit feedback, we perform a qubit reset after each time step. This ensures that no extraneous information (other than the intended feedback) propagates forward, allowing us to study the feedback coupling itself in a controlled manner. One can omit these resets to explore more complex scenarios in which correlations persist over time, but here our primary goal is to study feedback connections in isolation and thereby clarify their implications. For implementation, IBM superconducting systems have in particular demonstrated classical real-time feedforward operations [17], providing an opportunity to reduce the overall shot overhead for certain problems. We use the measurement strings obtained from a mid-circuit measurement as feedback for the next cycle. We can tune the feedback strength and the input strength through hyperparameters. We evaluate the memory performance by studying the short-term memory task and the predictive capabilities by predicting the Mackey-Glass system [18] and the expectation value of a spin in a chaotic Ising chain.
@@
Figure 1: QRC protocols: Each architecture is designed to process an input sequence ${s_{k}}$. The systems are initialized in the state $\rho_{0}$. A unitary operator $U(\cdot)$, constructed from parameterized Pauli rotations, governs the state evolution. Each $U(\cdot)$ represents the processing of one time step (cycle). (a) Graphical representation of a conventional restart-based QRC scheme. For each timestep, the architecture must be restarted and run from the beginning. This results in a quadratic runtime. (b) A rewind protocol that does not restart from the beginning, but only processes the last $\tau$ time steps. (c) A feedback controlled QRC protocol. Only one timestep is processed at a time, but the expectation value from the previous trial is stored in a buffer. (d) A mid-circuit measurement architecture that continuously monitors the quantum state. No restart is required. This architecture can be combined with weak measurements. (e) A feedback-driven QRC architecture with mid-circuit measurements. Feedback is incorporated by feeding the measurement strings $\textbf{m}_{k}$ back into the reservoir, a process that can be done at runtime.
@@ II Methods
In the following, we provide an overview of our approach in three parts: first we describe the proposed mid-circuit QRC architecture, then we present the training procedure, and finally we explain the testing routine.
@@ II-A Architecture
We propose a mid-circuit QRC architecture that incorporates feedback links, as shown in Fig. 1 (e). We start with a quantum system of $N$ qubits, initialized in the state $\rho_{0}^{N}$ which is chosen to be |$\ket{0}\bra{0}^{\otimes N}$ throughout this work unless otherwise mentioned. Given an input sequence ${s_{k}}$, the system evolution is governed by the unitary operator $U(s_{k},\textbf{m}_{k-1})$, where $s_{k}$ is the current time step of the input sequence and $\textbf{m}_{k-1}\in\{-1,1\}^{N}$ are previous measurement results of the cycle $s_{k-1}$. The input sequence is inserted into a two-qubit gate $R(a_{in}s_{k})$, while each previous measurement result $m_{k-1}^{j}$ is inserted into another two-qubit gate $R(a_{fb}m_{i}^{j})$. The structure of the $R$ gate is defined as follows:
@@
Here $\theta$ represents an angle parameter that determines the rotation applied in the two-qubit gate. Specifically, $\theta$ can take values from the set $\{a_{in}s_{k},a_{fb}m_{k-1}^{j}\}$, where $a_{in}$ and $a_{fb}$ are scaling parameters that control the input and feedback strength respectively. The structure of using $R$ gates along with the scaling parameters is commonly used in QRC models due a its compromise between being hardware efficient and generating substantial entanglement [9, 15, 19]. The full unitary $U(s_{k},\textbf{m}_{k-1})$ consists of an input $R$ gate, $N$ feedback $R$ gates, and a Haar-random unitary. The idea is illustrated for the case of $2$ qubits in Fig. 2. For the first time step $s_{0}$, a random vector $\textbf{m}_{-1}$ uniformly distributed from $\{-1,1\}^{N}$ is selected for time step 00. After unitary evolution, the system is measured projectively in the computational basis. We note that further research can also employ weak measurements, as seen in other QRC research [10, 15]. We reset our system to $\rho_{0}^{N}$ and feed the measurement string forward into the next unitary.
Figure 2: Minimal two-qubit unitary circuit. The circuit processes the input $s_{k}$ via a rotation $R(a_{\mathrm{in}}\,s_{k})$ and then incorporates feedback via rotations $R(a_{\mathrm{fb}}\,m_{k-1}^{0})$ and $R(a_{\mathrm{fb}}\,m_{k-1}^{1})$, where $\{m_{k-1}^{0},m_{k-1}^{1}\}$ are the previous measurement results. The structure of $R(\cdot)$ is explained in Eq. 1. A final Haar-random unitary $U_{\text{Haar}}$ is applied at the end.
@@ II-B Training
The task of QRC is to build a regression model that accurately approximates a target sequence $\{\hat{y}_{k}\}$ for a given input sequence $\{s_{k}\}$. In typical QRC mechanisms, we run the model for $l_{w}+l_{tr}+l_{ts}$ cycles, where $l_{w}$, $l_{tr}$, and $l_{ts}$ correspond to the lengths of the three main stages: washout, training, and testing respectively. As discussed in Section II-A, we start the first cycle with input $s_{0}$ and a randomly selected feedback vector. To remove the effects of this initialization, we discard the first $l_{w}$ measurement results (the washout phase). A regression model is then built using the next $l_{\mathrm{tr}}$ steps (the training phase). The linear regression model is built based on the reservoir outputs of the training phase. Specifically, we build the matrix $X_{tr}\in\mathcal{R}^{l_{tr}\times(N+1)}$, which is defined as:
@@
where $\langle Z\rangle_{k}^{n}$ is the expectation value of the Pauli $Z$ operator on the $n^{th}$ qubit at the $k^{th}$ time step.
We now create a linear model $w_{opt}$ such that $y_{tr}=X_{tr}w_{opt}$ closely resembles the target sequence $\hat{y}_{tr}$. We use the Moore-Penrose inverse [20, 21], also called the pseudoinverse, to compute $w_{opt}$:
@@ II-C Testing
We now test the trained model by using it to obtain a prediction vector of length $l_{ts}$. We create a matrix $X_{ts}\in\mathcal{R}^{l_{ts}\times(N+1)}$, which is defined as:
@@
The resulting prediction $y_{pred}$ is then compared to the target vector $\hat{y}_{ts}$ to evaluate the prediction performance of the model.
@@ III Experiments
In this section, we investigate the performance of our proposed reservoir computing model through three key experiments: measuring its short-term memory capacity, evaluating its predictive capabilities, and analyzing the echo state property.
@@ III-A Short-term memory capacity
To characterize the performance of remembering past inputs we use the short-term memory capacity [22]. Consider an input sequence $\{s_{k}\}$, where each $s_{k}$ is drawn independently from a uniform distribution on $[0,1]$. The goal is to retrieve values of the sequence, formulated as $\hat{y}_{k}=s_{k+\tau}$, where $\tau$ is zero or negative. We use the coefficient of determination, given by:
$R^{2}_{\tau}=\frac{\text{cov}^{2}(\hat{y}_{ts},y_{pred})}{\sigma^{2}(\hat{y}_{ts})\sigma^{2}(y_{pred})},$ (6)
as a figure of merit. When $R^{2}_{\tau}$ is close to 1, the model’s predictions closely match the actual targets, indicating minimal error. In contrast, when $R^{2}_{\tau}$ is close to 0, the predicted outputs show substantial deviation from the actual targets. Short-term memory capacity is then measured by summing $R^{2}$ for different time steps:
@@ III-B Predictive Capabilities
To evaluate the predictive capabilities, we test the model on a classical task and a quantum time series. For the study of classical dynamical systems, we choose the Mackey-Glass time series, a well-known benchmark that has been extensively used in the context of reservoir computing [23, 9, 5, 24]. The delay differential equation is defined as [18]:
@@
We set $\beta_{0}=0.2$, $\theta=1$, $n=10$, $\tau=17$ and $\gamma=0.1$. To generate the time series, we numerically integrate the delay differential equation using a forward Euler scheme with a time step of $\Delta t=1$. The initial values are set to a constant for the duration of the delay, and the trajectory is normalized to the interval $[0,1]$.
The quantum time series is constructed from the expectation value of the $\sigma^{z}$ operator of the middle spin in a one-dimensional Ising chain governed by the Hamiltonian:
@@
We simulate the system using this Hamiltonian with parameters set to $N=5$, $J=1$, $h_{x}=1.05$, and $h_{z}=-0.5$. This ensures that the quantum Ising model becomes non-integrable and exhibits chaotic behavior [25, 26]. At each time step, the expectation value $\langle Z\rangle_{t}^{2}$ of the central spin is measured, generating the time series. The time evolution is discretized with a time step of $\Delta t=0.005$.
For all prediction tasks, the goal is to predict steps in the future, similarly formulated with the target sequence as $\hat{y}_{k}=s_{k+\tau}$, where $\tau$ is positive. We evaluate the performance of the models with the norm mean squared error, which is defined as:
@@
The NMSE measures the size of the error relative to the scale (magnitude) of the true data. A small error indicates good model performance, while large values indicate a large deviation from the target sequence.
@@ III-C Analysis of Echo State Property
As mentioned in Section I, one of the important aspects of a reservoir computing model is the fading memory property. In the classical setting of echo state network (ESN) reservoir computing models, this can be implemented by ensuring that the ESN contains the echo state property (ESP) [27], which ensures that the internal state of the model becomes asymptotically dependent only on the input history, regardless of its initial conditions. However, in quantum reservoir models, the fading memory property is usually generated by including some (preferably controlled) dissipation in the model [28, 29]. Recently, there have been proposals of modified versions of the ESP that are specific to quantum reservoirs [30]. In this paper, we describe and investigate a version of the ESP that applies across any reservoir computing model.
As a first step in investigating the role of the ESP in quantum reservoir computing models, we perform an analysis of the asymptotic dependence of the reservoir state of several reservoir computing models (including the one proposed in this paper) on their initial internal states. To do this, we study how the internal state of the reservoir evolves over time under different random initializations. We can define the internal state of the reservoir the same way as is defined in traditional reservoir computing models, i.e., $x_{t}\in\mathcal{R}^{M}$ at time $t$ to be
@@
where $M$ is the reservoir dimension, $F$ is the fixed reservoir function, and $s_{t}$ is the time series at time $t$ [27, 14]. [^1] In our architecture, $x_{t}$ = $[\langle Z\rangle^{1}_{t},\langle Z\rangle^{2}_{t},\cdots,\langle Z\rangle^{N}_{t}]$.. For each model, we use input sequences drawn from a uniform distribution on $[0,1]$ (as described in Section III-A) over 5 random initializations of the model. For each run, we extract the first component of the internal state of the reservoir, $x_{t}^{1}$, at time $t$ and track its evolution over time across the different runs. To quantify the asymptotic behavior of these internal states, we compute the average difference of this component across runs. We perform this analysis for a classical minimal ESN model [31], a feedback-driven QRC protocol [9], a QRC scheme with a mid-circuit measurements architecture [15], and the model proposed in this paper.